prettygirl15:

A golf ball is thrown off of the roof of a building. It’s height h(t) after t seconds is modeled by the function ℎ ( 푡 ) = − 5 푡 2 + 4 푡 + 18 What is the initial height of the golf ball?

xXAikoXx:

So basically, this is solve for h instead of 𝑥 h= -5 2+4+18 Solve for h. −52+4+18 =−48+18 =−30 =−30

xXAikoXx:

h = -30

prettygirl15:

@xxaikoxx wrote:
h = -30
thx

dude:

Hmm your question seems to have been mixed with other characters Is the equation given \(h(t)=-5t^2+4t+18\)? If you want to find initial height, set t equal to 0

@xxaikoxx wrote:
h = -30
Careful! The question asks for height. If h is -30 then this means the ball was thrown below ground which does not match the question